OpenStudy (anonymous):

How many solutions can be found for the equation –4x – 11 = 2(x – 3x) + 13? None One Two Infinitely many

OpenStudy (alexandervonhumboldt2):

solve: -4x-11=2*(x-3x)+13 use distributive property: -4x-11=2x-6x+13

sammixboo (sammixboo):

Well we can start this off by simplifying the equation. Do you know how to simplify it some?

OpenStudy (alexandervonhumboldt2):

now simplify the 2x-6x part. -4x-11=-4x+13 0=26 which is of course false

sammixboo (sammixboo):

It isn't 26

OpenStudy (alexandervonhumboldt2):

oops 24 mistyped

OpenStudy (anonymous):

so is it the first second third or fourth?

OpenStudy (anonymous):

help!!

sammixboo (sammixboo):

Well if it is false, then that means there is no answer

OpenStudy (anonymous):

Please don't give out answers. :(

OpenStudy (anonymous):

well they are helpfull

OpenStudy (anonymous):

It does not matter. Ask your teacher.

OpenStudy (anonymous):

lol go away!

OpenStudy (anonymous):

Reporting.

sammixboo (sammixboo):

Giving direct answers isn't helping you learn the process of how to solve these problems.